What are the monthly payments (principal and interest) on a 15-year home mortgage for an $160,000 loan when interest rates are fixed at 7 percent? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASEENG 03:45 30-07-2020 F Q 99+ 32 Q869 . L M N P Q R S o T 851 852 853 854 BA II PLUS 855 set C/Y = P/Y = 12 -160000 856 857 85